# Benjamin S. Manning MIT John J. Horton

### v3

@@ Abstract
Useful social science theories predict behavior across settings. However, applying a theory to make predictions in new settings is challenging: rarely can it be done without ad hoc modifications to account for setting-specific factors. We argue that AI agents put in simulations of those novel settings offer an alternative for applying theory, requiring minimal or no modifications. We present an approach for building such “general” agents that use theory-grounded natural language instructions, existing empirical data, and knowledge acquired by the underlying AI during training. To demonstrate the approach in settings where no data from that data-generating process exists—as is often the case in applied prediction problems—we design a highly heterogeneous population of 883,320 novel games. AI agents are constructed using human data from a small set of conceptually related, but structurally distinct “seed” games. In preregistered experiments, on average, agents predict human play better than (i) game-theoretic equilibria and (ii) out-of-the-box agents in a random sample of 1,500 games from the population. For a small set of separate novel games, these simulations predict responses from a new sample of human subjects better even than the most plausibly relevant published human data.
@@ 1 Introduction
A general, low-cost method for accurately simulating human behavior with AI agents would have wide application in the social sciences (Charness et al., 2025; Jackson et al., 2025). Recognizing this potential, a growing literature explores whether large language models (LLMs) can simulate human responses in various settings.111(Horton, 2023; Argyle et al., 2022; Aher et al., 2023; Mei et al., 2024; Wang et al., 2025; Park et al., 2024, 2023; Chang et al., 2024; Binz and Schulz, 2023, 2024; Cerina and Duch, 2025; Manning et al., 2024; Capra et al., 2024; Shah et al., 2025; Anthis et al., 2025; Hansen et al., 2024; Broska et al., 2025; Bybee, 2025; Kim et al., 2024) Across dozens of experiments, samples of these agents respond with remarkable similarity to humans—even when simulating novel studies that did not appear in the underlying LLM’s training corpus (Hewitt et al., 2024; Binz et al., 2024; Li et al., 2024; Suh et al., 2025; Tranchero et al., 2024). Yet within this literature, others find settings where the very same models are poor human proxies.222(Santurkar et al., 2023; Atari et al., 2023; Cheng et al., 2023; Gao et al., 2024; Gui and Toubia, 2023) This inconsistency poses a challenge for AI simulations as robust and credible predictive models—particularly in settings where no prior human data exists. The core challenge is not simply achieving a close match between AI and human responses in one setting, but building agents that will generalize reliably.
@@
A natural starting point is improving the instructions given to agents. These instructions, or “prompts,” are written descriptions given to the LLM specifying who it is, what it believes, or how it should behave and reason. Such second-person instructions (e.g., “You respond as a type-X person”) can profoundly affect output distributions because advanced LLMs have been explicitly fine-tuned to follow instructions (Bai et al., 2022; Ouyang et al., 2022). With an appropriate prompt (which can now be massive), many highly capable models can perform complex reasoning and mathematical tasks at levels sometimes better even than those of highly skilled humans. [^3] To date, Google’s Gemini 1.5 can accommodate 10 million tokens (Gemini Team, 2024). This is roughly equivalent to 15,000 pages, or about 20 copies of the Handbook of Experimental Economics (Kagel and Roth, 1995).
@@
Despite the existence of these powerful, steerable models, constructing agents whose behavior is similar to that of real humans in a wide variety of settings is nontrivial—even when human data are available to guide the search. The set of possible prompts is vast, ranging from simple combinations of social or demographic traits to complex programmatic instructions related to how humans make decisions (Zhu et al., 2025; Xie et al., 2025). As in other machine-learning applications, the challenge is not only to avoid underfitting but also to guard against overfitting. By iterating through enough prompts, one can almost always find some arbitrary prompt that shifts the LLM’s responses to closely match a given human distribution. For example, an LLM instructed “you randomly offer between \$6 and \$9” may perfectly reproduce a distribution of human responses in a \$20 dictator game, but such a prompt would be nonsensical for a \$5 dictator game. In contrast, a prompt grounded in the underlying behavioral drivers—e.g., “you are self-interested but fair”—can perform well in-sample and plausibly extend to a range of allocation games. Standard data-driven approaches, such as a train-test split within a single dataset, cannot reliably distinguish between these two cases; the latter appears better only when tested in truly new settings. If the goal is to predict behavior in settings with no prior human data, how should researchers construct prompts?
@@
The approach uses two kinds of data: i) training data—existing human data used to optimize AI simulations, and ii) validation or test data—existing human data related to but distinct from the training data, used to test whether the optimized agents generalize. The ultimate goal is to produce agents that can more accurately predict human behavior in novel target settings where no prior human data exist, but that lie in the same broad domain as the training and validation settings.
The first step of the approach is to limit the “space” of prompts to a subset motivated by some economic theory or causal mechanism relevant to the novel setting of interest. This theory-grounding is analogous to constraining the functional form of the hypothesis class in machine learning. Continuing with the dictator game example, where social preferences likely determine behavior, one might choose the candidate set of prompts characterized by known drivers of the relevant preferences (e.g., “You are {level}” for all levels in {self-interested but fair, altruistic, selfish}).
The second step is to optimize over this set to best match the human training data. [^4] There is an active literature on optimizing prompts (Khattab et al., 2024). All candidate prompts are put in simulations of the settings that produced the training data (e.g., a \$20 dictator game). Optimization effectively filters these candidate prompts down to a subset whose simulated responses are closest to the human training distribution. We employ two optimization methods: (i) a selection method that identifies the optimal mixture of prompts from the candidate set (Xie et al., 2025; Leng et al., 2024; Bui et al., 2025), and (ii) a construction method that optimizes numerical parameters embedded directly in the prompts. Both methods produce a final set of prompts. If the set achieves good in-sample fit, we have reason to believe they will generalize because they are also grounded in a relevant theory. Poor in-sample fit suggests a mismatch between the theory and the training setting or an inadequate operationalization of that theory. In either case, we revise the candidate prompts and re-optimize.
@@
We illustrate this approach and provide evidence of its efficacy with training and validation data drawn from experiments in the behavioral economics literature applied to AI agents. All simulations use GPT-4o with temperature set to 1, though the approach is agnostic to the choice of model and hyperparameters. [^5] Gpt-4o, among the most widely used LLMs, and the temperature setting are defaults for the software used to run our simulations (Horton and Horton, 2024). We first apply the selection method to Arad and Rubinstein (2012)’s 11-20 money request game, where participants request an amount and receive a bonus if they choose exactly one less than their opponent. We use only the original dataset from the paper, which contained fewer than 200 observations. This setting is appealing to study because optimal play depends not only on the focal agent’s capabilities, but also on their beliefs about others’ strategic reasoning. Endowing AI agents with distinct prompts corresponding to varying degrees of strategic reasoning (the theoretical focus of Arad and Rubinstein) produces a mixture that closely matches the original human data.
When we validate these optimized samples on distinct variants of the 11-20 money request game, they are better predictors of human responses than baseline AI agents with no additional instructions. By contrast, scientifically meaningless or “atheoretical” AI agents derived from historical figures, pseudo-scientific Myers-Briggs personality types, and those instructed to select particular numbers can sometimes match human distributions in one variant of the game but fail to generalize across others.
@@
What statistical guarantees does this approach afford? Without a correctly specified causal model, no statistical procedure can guarantee performance in arbitrary new environments (Pearl, 2009). Formal guarantees with existing data, like those required for prediction-powered inference (Angelopoulos et al., 2023) and other related methods (Egami et al., 2023; Hardy et al., 2025), would require a strictly firewalled validation set: data never used in the construction of the underlying LLM (Ludwig et al., 2024; Sarkar and Vafa, 2024; Mullainathan and Spiess, 2017; Modarressi et al., 2025). This is a tall order impossible to meet in practice, even with existing public weight LLMs, never mind private models. However, what we can guarantee is prediction performance over a pre-committed family of settings.
@@
We explore this approach to inference by constructing a population of 883,320 novel strategic games. These were inspired by the original 11-20 game but were modified along several dimensions, such as support of possible choices, the size and nature of the bonus, the manner in which money is allocated, etc. The population of games is the full factorial combination of these dimensions. The resulting games differed substantively from each other; in some cases, they were cooperative, others adversarial, many had a dominant strategy equilibrium, while some had no dominant strategies at all. The vast majority are surely not in any LLM’s training corpus to date. From the population, we sample 1,500 unique games and have 4,500 human subjects each play a game in a final preregistered experiment. We take the theoretically-grounded level-$k$ agents—agents constructed months before the novel set of 883,320 games existed—and evaluate their ability to predict the human responses.
@@
In a final preregistered experiment, these theory-grounded agents are far better predictors of the human data than the baseline AI off-the-shelf. In the average game, they assign $3.41$ times more probability to the actions actually taken by human subjects. The optimized agents are similarly effective ($2.44$ times) compared to the unique Harsanyi-Selten-selected equilibria (Harsanyi and Selten, 1988) calculated for each of the 1,500 games. Because the games were randomly sampled, the corresponding confidence intervals over the relative predictive power are externally valid for the much larger population.
@@
The goal of AI simulations in this paper is to harness two extensive sources of information to create better predictive agents: i) well-established theoretical models from the social sciences, and ii) the immense knowledge about human behavior that LLMs have implicitly learned during training (Lindsey et al., 2025; Ameisen et al., 2025). In a sense, AI agents are a general vessel through which theory can be flexibly applied to any setting. Our approach aims to generate such agents by conducting a structured trial run across various settings, building evidence that theoretically-grounded prompts generalize effectively to similar yet distinct environments. These principles can be applied to any domain where relevant human data exists—even those with multiple interacting agents. One can imagine iteratively identifying novel scenarios where predictions are needed, finding the closest relevant existing data, and continuously optimizing and calibrating agents as needed. [^7] LLMs could automatically construct candidate prompts as input to such feedback loops (Xie et al., 2025). Our main contribution is to systematize these ideas for constructing agents and then provide empirical evidence that they can substantially improve the predictive power of AI simulations in new settings.
@@
The remainder of this paper is organized as follows. Section 2 begins with a concrete example of identifying generalizable relationships and constructing prompts that better predict human subjects in new settings. We then describe the idea more generally and specify the optimization methods. Section 3 illustrates the approach and demonstrates its efficacy empirically with several experiments. Section 4 demonstrates how agents can be used in a pre-committed setting to provide externally valid statistical estimates of predictive accuracy at scale. The paper concludes in Section 5.
@@ 2 Building prompts that generalize
This section explains the approach for constructing AI agents that better approximate human response distributions in new environments. We begin with an extended example of predicting behavior in a novel public goods game where no previous human data exists. We then discuss the importance of validating across distinct data-generating processes and grounding candidate prompts in relevant social science theories. The section concludes with a summary of the approach and a brief description of the optimization methods used to construct agents.
@@ 2.1 A motivating example
Suppose we want to predict how people will share resources in a novel public goods game (as they do in Alsobay et al. (2025)). In this new game, for which we have no previous data, three participants will be endowed with \$5 and can choose to contribute any portion of their endowment, which will be multiplied by 3 and then divided equally among all participants. While we do not have any prior public-goods game data, we do have human data from a related \$20 dictator game. It is related in the sense that one might reasonably expect some generalizable features of human choice to affect allocations in both games. The observed human offers from the dictator game are {6,6,7,7,8,8,9,9}.
@@
Suppose that, prompted with the dictator game instructions, the LLM produces a response distribution {3,3,3,3,3,4,4,4}. Although well within the allowable offers, this distribution clearly diverges from the observed human data ({6,6,7,7,8,8,9,9}) and leaves us with little hope that it could effectively predict responses to the novel public goods game. Thus, even though the model may capture some general aspects of human behavior as a baseline, i.e., a tendency to offer nontrivial amounts (Henrich et al., 2001), it does so imperfectly.
One might think this problem could be addressed by first randomly splitting the human sample of dictator game offers into training and testing sets. Then identifying the best-performing prompts on the training set, and validating their performance on the test set (Mullainathan and Spiess, 2017; Ludwig et al., 2024). Indeed, an LLM instructed “You randomly choose numbers between 6 and 9” could reasonably predict both training and testing sets for any split of the human data better than the baseline LLM with no persona. However, such an approach will almost certainly fail to generalize beyond the specific data-generating process that produced the dictator game offers. It has, in effect, still overfit. Rather than overfitting to the training data, it has overfit to the whole data-generating process.
@@ 2.2 Identifying generalizable behavioral relationships
This validation process becomes even more robust when multiple distinct but theoretically-related datasets are available. Imagine dictator-game responses from \$1, \$5, \$10, \$20, \$50, and \$100 games, all exhibiting offers slightly below half. Optimizing and validating across these multiple settings reduces the risk of overfitting. With each new training data-generating process, it is less likely that an arbitrary prompt will generalize in-sample. [^8] We do not offer empirical examples of optimizing over multiple training settings in the main text (only single settings in Section 3). Appendix A provides such analyses, including an additional preregistered experiment. Thus, if the underlying relationship governing these settings also holds for a novel target setting, the validated prompts should robustly predict behavior there as well.
Yet, even with an effective validation method, a fundamental practical challenge remains: identifying the initial set of candidate prompts. While our motivating example made this step appear straightforward, selecting plausible prompts in more complex settings can be far less obvious. In a perfect world, an optimization procedure would take in a massive set of prompts and filter them down to a smaller set that generalizes. However, this does not yet work in practice because many different prompts can be used to generate the same behavior in a given setting. For example, the number of possible natural language prompts that get an LLM to respond with numbers between 6 and 9 is enormous. If we were to start from a large pool of prompts, then the optimization procedure might endlessly produce sets of prompts that fit in-sample but fail on the validation set. For now, it is essential to ground the candidate prompts with a principled starting point, which we argue, and later demonstrate empirically, that economic and behavioral theories offer.

### v4

@@ Abstract
Useful social science theories predict behavior across settings. However, applying a theory to make predictions in new settings is challenging: rarely can it be done without ad hoc modifications to account for setting-specific factors. We argue that AI agents put in simulations of those novel settings offer an alternative for applying theory, requiring minimal or no modifications. We present an approach for building such “general” agents that use theory-grounded natural language instructions, existing empirical data, and knowledge acquired by the underlying AI during training. To demonstrate the approach in settings where no data from that data-generating process exists—as is often the case in applied prediction problems—we design a heterogeneous population of 883,320 novel games. AI agents are constructed using human data from a small set of conceptually related but structurally distinct “seed” games. In preregistered experiments, on average, agents predict initial human play in a random sample of 1,500 games from the population better than (i) a cognitive hierarchy model, (ii) game-theoretic equilibria, and (iii) out-of-the-box agents. For a small set of separate novel games, these simulations predict responses from a new sample of human subjects better even than the most plausibly relevant published human data.
@@ 1 Introduction
A general, low-cost method for accurately simulating human behavior with AI agents would have wide application in the social sciences (Charness et al., 2025; Jackson et al., 2025; Anthis et al., 2025). Recognizing this potential, a growing literature explores whether large language models (LLMs) can simulate human responses in various settings.111(Argyle et al., 2022; Aher et al., 2023; Binz and Schulz, 2023; Brand et al., 2023; Park et al., 2023; Mei et al., 2024; Park et al., 2024; Chang et al., 2024; Manning et al., 2024; Capra et al., 2024; Hansen et al., 2024; Kim et al., 2024; Wang et al., 2025; Cerina and Duch, 2025; Shah et al., 2025; Broska et al., 2025; Bybee, 2025; Fish et al., 2025) Across dozens of experiments, samples of these agents respond with remarkable similarity to humans—even when simulating novel studies that did not appear in the underlying LLM’s training corpus (Hewitt et al., 2024; Binz et al., 2024; Li et al., 2024; Tranchero et al., 2024; Suh et al., 2025). Yet within this literature, others find settings where the very same models are poor human proxies.222(Santurkar et al., 2023; Atari et al., 2023; Cheng et al., 2023; Gui and Toubia, 2023; Gao et al., 2024) This inconsistency poses a challenge for AI simulations as robust and credible predictive models—particularly in settings where no prior human data exists. The core challenge is not simply achieving a close match between AI and human responses in one setting, but building agents that will generalize reliably.
@@
A natural starting point is improving the instructions given to agents. These instructions, or “prompts,” are written descriptions given to the LLM specifying who it is, what it believes, or how it should behave and reason (Horton, 2023). Such second-person instructions (e.g., “You respond as type-X person”) can profoundly affect output distributions because advanced LLMs have been explicitly fine-tuned to follow instructions (Bai et al., 2022; Ouyang et al., 2022). With an appropriate prompt (which can be massive), [^3] To date, Google’s Gemini 1.5 can accommodate 10 million tokens (Gemini Team, 2024). This is roughly equivalent to 15,000 pages, or about 20 copies of the Handbook of Experimental Economics (Kagel and Roth, 1995). highly capable models can perform complex reasoning and mathematical tasks at levels sometimes better even than those of highly skilled humans.
@@
Despite the existence of these powerful, steerable models, constructing agents whose behavior is similar to that of real humans in a wide variety of settings is nontrivial—even when human data are available to guide the search. The set of possible prompts is vast, ranging from simple combinations of social or demographic traits to complex programmatic instructions related to how humans make decisions (Zhu et al., 2025b; Xie et al., 2025). As in other machine learning applications, the challenge is not only to avoid underfitting but also to guard against overfitting. By iterating through enough prompts, one can almost always find some arbitrary prompt that shifts the LLM’s responses to closely match a given human distribution. For example, an LLM instructed “you randomly offer between \$6 and \$9” may perfectly reproduce a distribution of human responses in a \$20 dictator game, but such a prompt would be nonsensical for a \$5 dictator game. In contrast, a prompt grounded in the underlying behavioral drivers—e.g., “you are self-interested but fair”—can perform well in-sample and plausibly extend to a range of allocation games. Standard data-driven approaches, such as a train-test split within a single dataset, cannot reliably distinguish between these two cases; the latter appears better only when tested in truly new settings. If the goal is to predict behavior in settings with no prior human data, how should researchers construct prompts?
@@
The approach uses two kinds of data: (i) training data—existing human data used to optimize AI simulations, and (ii) validation or test data—existing human data related to but distinct from the training data, used to test whether the optimized agents generalize. The ultimate goal is to produce agents that can more accurately predict human behavior in novel target settings where no prior human data exist, but that lie in the same broad domain as the training and validation settings.
The first step of the approach is to limit the “space” of prompts to a subset motivated by some economic theory or causal mechanism relevant to the novel setting of interest. This theory-grounding is analogous to constraining the functional form of the hypothesis class in machine learning. Continuing with the dictator game example, where social preferences likely determine behavior, one might choose the candidate set of prompts characterized by known drivers of the relevant preferences (e.g., “You are {level}” for all levels $\in$ {self-interested but fair, altruistic, selfish}).
The second step is to optimize over this set to best match the human training data. [^4] There is an active literature on optimizing prompts (Khattab et al., 2024). All candidate prompts are put in simulations of the settings that produced the training data (e.g., a \$20 dictator game). Optimization effectively filters these candidate prompts down to a final subset whose simulated responses are closest to the human training distribution. We employ two optimization methods: (i) a selection method that identifies the optimal mixture of prompts from the candidate set (Leng et al., 2024; Xie et al., 2025; Bui et al., 2025), and (ii) a construction method that optimizes numerical parameters embedded directly in the prompts. If the final optimized set achieves a good in-sample fit, we have reason to believe they will generalize because they are also grounded in a relevant theory. Poor in-sample fit suggests a mismatch between the theory and the training setting or an inadequate operationalization of that theory. In this case, we revise the candidate prompts and re-optimize.
@@
We illustrate this approach and provide evidence of its efficacy with training and validation data drawn from experiments in the behavioral economics literature. All simulations use Gpt-4o with the temperature set to 1, though the approach is agnostic to the choice of model and hyperparameters. [^5] Gpt-4o, among the most widely used LLMs, and the temperature setting are defaults for the software used to run our simulations (Horton and Horton, 2024). We first apply the selection method to Arad and Rubinstein (2012)’s 11-20 money request game, where participants request an amount and receive a bonus if they choose exactly one less than their opponent. We use only the original dataset from the paper, which contained fewer than 200 observations. This setting is appealing to study because optimal play depends not only on the focal agent’s capabilities, but also on their beliefs about how others will reason. Endowing AI agents with distinct prompts corresponding to varying degrees of strategic reasoning (the theoretical focus of Arad and Rubinstein) produces a mixture that closely matches the original human data.
When we validate these optimized samples on distinct variants of the 11-20 money request game, they are better predictors of initial human play than baseline AI agents with no additional instructions. By contrast, scientifically meaningless or “atheoretical” AI agents derived from historical figures, pseudo-scientific Myers-Briggs personality types, and those instructed to select particular numbers can sometimes match human distributions in one variant of the game but fail to generalize across others.
@@
What statistical guarantees does this approach afford? Without a correctly specified causal model, no statistical procedure can guarantee performance in arbitrary new environments (Pearl, 2009). Formal guarantees with existing data, like those required for prediction-powered inference (Angelopoulos et al., 2023) and other related methods (Egami et al., 2023; Hardy et al., 2025), would require a strictly firewalled validation set: data never used in the construction of the underlying LLM (Ludwig et al., 2025; Sarkar and Vafa, 2024; Mullainathan and Spiess, 2017; Modarressi et al., 2025). This is a tall order impossible to meet in practice, even with existing public weight LLMs, never mind private models. However, what we can guarantee is prediction performance over a pre-committed family of settings.
@@
We explore this approach to inference by constructing a population of 883,320 novel strategic games. These were inspired by the original 11-20 game but were modified along several dimensions, such as support of possible choices, the size and nature of the bonus, the manner in which money is allocated, etc. The population of games is the full factorial combination of these dimensions. The resulting games differed substantively from each other; in some cases, they were cooperative, others zero-sum, many had a dominant strategy equilibrium, while some had no dominant strategies at all. The vast majority are surely not in any LLM’s training corpus to date. From the population, we sample 1,500 unique games and have 4,500 human subjects each play a game in a final preregistered experiment. We take the theoretically-grounded level-$k$ agents—agents constructed months before the novel set of 883,320 games existed—and evaluate their ability to predict the human responses.
@@
The theory-grounded agents are far better predictors of initial human play than the baseline AI off-the-shelf. In the average game, they assign $3.41$ times more probability to the actions actually taken by human subjects. The optimized agents are similarly effective compared to two well-established theoretical benchmarks: (i) a symmetric Nash equilibrium calculated for each of the 1,500 games ($2.44$ times more probability), and (ii) the game-specific predictions from the cognitive hierarchy model of Camerer et al. (2004) ($3.02$ times more probability). Because the games were randomly sampled, the corresponding confidence intervals over the relative predictive power are externally valid for the much larger population.
@@
The goal of AI simulations in this paper is to harness two extensive sources of information to create better predictive agents: (i) well-established theoretical models from the social sciences, and (ii) the immense knowledge about human behavior that LLMs have implicitly learned during training (Lindsey et al., 2025; Ameisen et al., 2025). In a sense, AI agents are a general vessel through which theory can be flexibly applied to any setting. Our approach aims to generate such agents by conducting a structured trial run across various settings, building evidence that theoretically-grounded prompts generalize effectively to similar yet distinct environments. These principles can be applied to any domain where relevant human data exists—even those with multiple interacting agents (Manning et al., 2024; Qian et al., 2025). One can imagine iteratively identifying novel scenarios where predictions are needed, finding the closest relevant existing data, and continuously optimizing and calibrating agents as needed. [^7] LLMs could automatically construct candidate prompts as input to such feedback loops (Xie et al., 2025). Our main contribution is to systematize these ideas for constructing agents and then provide empirical evidence that they can substantially improve the predictive power of AI simulations in new settings.
@@
The remainder of this paper is organized as follows. Section 2 begins with a concrete example of identifying generalizable relationships and constructing prompts that better predict human subjects in new settings. We then describe the approach more generally and specify the optimization methods. Section 3 illustrates the approach and demonstrates its efficacy empirically with several experiments. Section 4 demonstrates how agents can be used in a pre-committed setting to provide externally valid statistical estimates of predictive accuracy at scale. The paper concludes in Section 5.
@@ 2 Building prompts that generalize
This section explains the approach for constructing AI agents that better approximate human response distributions in new settings. We begin with an extended example of predicting behavior in a novel public goods game where no previous human data exists. We then discuss the importance of validating across distinct data-generating processes and grounding candidate prompts in relevant social science theories. The section concludes with a summary of the approach and a brief description of the optimization methods used to construct agents.
@@ 2.1 A motivating example
Suppose we want to predict how people will share resources in a novel public goods game (as they do in Alsobay et al. (2025)). In this new game, for which we have no previous data, three participants will be endowed with \$5 and can choose to contribute any portion of their endowment, which will be multiplied by 3 and then divided equally among all participants. While we do not have any prior public goods game data, we do have human data from a related \$20 dictator game. It is related in the sense that one might reasonably expect some generalizable features of human choice to affect allocations in both games. The observed human offers from the dictator game are {6,6,7,7,8,8,9,9}.
@@
Suppose that, when prompted with the dictator game instructions, the LLM produces a response distribution {3,3,3,3,3,4,4,4}. Although well within the allowable offers, this distribution clearly diverges from the observed human data ({6,6,7,7,8,8,9,9}) and leaves us with little hope that it could effectively predict responses to the novel public goods game. Thus, even though the model may capture some general aspects of human behavior as a baseline, i.e., a tendency to offer nontrivial amounts (Henrich et al., 2001), it does so imperfectly.
One might think this problem could be addressed by first randomly splitting the human sample of dictator game offers into training and testing sets. And then identifying the best-performing prompts on the training set, and validating their performance on the test set (Mullainathan and Spiess, 2017; Ludwig et al., 2025). Indeed, an LLM instructed “You randomly choose numbers between 6 and 9” could reasonably predict both training and testing sets for any split of the human data better than the baseline LLM with no persona. However, such an approach will almost certainly fail to generalize beyond the specific data-generating process that produced the dictator game offers. It has, in effect, still overfit. Rather than overfitting to the training data, it has overfit to the whole data-generating process.
@@ 2.2 Identifying generalizable behavioral relationships
This validation process becomes even more robust when multiple distinct but theoretically-related datasets are available. Imagine dictator-game responses from \$1, \$5, \$10, \$20, \$50, and \$100 games, all exhibiting offers slightly below half. Optimizing over multiple settings simultaneously further reduces the risk of overfitting. With each new training data-generating process, it is less likely that an arbitrary prompt will generalize in-sample. [^8] We do not offer empirical examples of optimizing over multiple training settings in the main text (only single settings in Section 3). Appendix A provides such analyses, including an additional preregistered experiment. Thus, if the underlying relationship governing these settings also holds for a novel target setting, the validated prompts should robustly predict behavior there as well.
Yet, even with an effective validation method, a fundamental practical challenge remains: identifying the initial set of candidate prompts. While our motivating example made this step appear straightforward, selecting plausible prompts in more complex settings can be far less obvious. In a perfect world, an optimization procedure would take in a massive set of prompts and filter them down to a smaller set that generalizes. However, this does not yet work in practice because many different prompts can be used to generate the same behavior in a given setting. For example, the number of possible natural language prompts that get an LLM to respond with numbers between 6 and 9 is enormous. If we were to start from a large pool of prompts, then the optimization procedure might endlessly produce sets of prompts that fit in-sample but fail on the validation set. For now, it is essential to ground the candidate prompts with a principled starting point, which—we argue and later demonstrate empirically—economic and behavioral theories offer.